OpenStudy (cydney_morgan):

Is the correct answer 10 percent? As one animal eats another animal, about ______ percent of the energy is lost as heat. five ten twenty five ninety

OpenStudy (cydney_morgan):

@Abhisar Am I correct on this one? Or is it 25?

OpenStudy (chrisisdope1):

d

OpenStudy (chrisisdope1):

Its 90% of what is consumed eventually dissipates as heat from chemical reactions such as those that happen when muscles work and those that produce the basic body temperature even when asleep. The reactions that keep you alive all lose some energy to heat.

OpenStudy (abhisar):

10% is transferred

OpenStudy (abhisar):

So 90% is lost !

OpenStudy (cydney_morgan):

So I was kind of half correct?

OpenStudy (chrisisdope1):

10% is still alive!!!

OpenStudy (abhisar):

kinda..but there is nothing like half correct..it's either correct ot incorrect :D

OpenStudy (cydney_morgan):

But is the question asking how much energy is transported through each organism or is it asking how much is transferred to heat in general?

OpenStudy (cydney_morgan):

Because then it could be 90% or 10%.

OpenStudy (chrisisdope1):

I Just Explained It READ

OpenStudy (cydney_morgan):

Sorry :c

OpenStudy (abhisar):

It's asking about how much energy is lost in between transfer from one tropic level to another.

OpenStudy (abhisar):



OpenStudy (abhisar):

So you can see that in every transaction between two tropic levels, 90% of energy is lost and only 10% is transferred. Getting it ?

OpenStudy (cydney_morgan):

OOOH I get it! But I thought the Law of Conservation of Energy said Energy cannot be created nor destroyed, or lost? Am I completely wrong here? Sorry! >.<

OpenStudy (cydney_morgan):

I'm confusing myself I am so sorry for making this complicated.

OpenStudy (abhisar):

The energy lost doesn't means it is destroyed. It just escaped from a particular system (living being) to other(Environment)

OpenStudy (chrisisdope1):

Its Okay I Didnt Get It At First To :)

OpenStudy (abhisar):

\(\color{blue}{\text{Originally Posted by}}\) @chrisisdope1 \(\cancel{\bcancel{\sf It~Can~Be~Destroyed}}\) \(\color{blue}{\text{End of Quote}}\)

OpenStudy (chrisisdope1):

I meant Cant w.e

OpenStudy (abhisar):

`Energy can't be destroyed neither be created !` It only changes one form to another.

OpenStudy (cydney_morgan):

Oh ok I see. So the answer for my question in fact is not 10%,it is 90%?

OpenStudy (abhisar):

yes !

OpenStudy (chrisisdope1):

I Said Its 90% 3 TIMES!!!

OpenStudy (cydney_morgan):

I am sorry! I am not here to cheat! I am here to learn and I wanted to be sure I actually understood everything!

OpenStudy (abhisar):

it's ok...no need for sorry ! we are here to make u learn only.. \(\Huge{\overset{\frown}{\normalsize \left( \begin{matrix} \Large\cdot \quad \cdot\\ \cdot\\ \huge \smile \end{matrix} \right)}}\normalsize \\ \;/\quad \;\;\quad \backslash\)